# RealEngine: Simulating Autonomous Driving in Realistic Context

### v1

@@ A.3 LiDAR reconstruction
GS-LiDAR jiang2025gslidar 0.315 6.930 21.25 0.0903 20.88
We use GS-LiDAR jiang2025gslidar to reconstruct the LiDAR point clouds. GS-LiDAR jiang2025gslidar introduces a novel panoramic rendering technique with explicit ray-splat intersection, guided by panoramic LiDAR supervision. This method employs 2D Gaussian primitives huang20242d with periodic vibration characteristics chen2023periodic , allowing for precise geometric reconstruction of both static and dynamic elements in driving scenes.
The LiDAR point clouds in nuPlan nuplan is generated by merging point clouds from five LiDAR sensors, with non-uniform angular distributions across different scan lines. However, GS-LiDAR assumes that the LiDAR point cloud originate from a single laser scanner with uniformly spaced scan lines. To address this discrepancy, we reproject the merged point cloud onto the top LiDAR sensor to obtain the range map, selecting the point with the smaller depth when multiple points project to the same pixel. Additionally, we discard the non-uniform scan lines at both ends while retaining the central region, where the scan lines are evenly distributed.
As shown in Figure 10, this processing method results in some loss of LiDAR information. However, since the driving agent Chitta2023PAMI ; liao2024diffusiondrive converts the LiDAR point cloud into a 2-bin histogram over a 2D BEV grid with relatively low resolution, the missing information has a minimal impact on the driving agent. We evaluate the reconstruction performance on the same 6 sequences as the camera image reconstruction in Table 6.
Figure 10: LiDAR reconstruction. Although reprojection may lead to some loss of LiDAR information, its impact on the low-resolution histogram used by the agent is minimal. Meanwhile, GS-LiDAR jiang2025gslidar achieves high-quality reconstruction of the reprojected LiDAR data.
Appendix B Consistency of driving behaviors
We replicate the same real-world evaluation scenarios from the Navsim Dauner2024NEURIPS benchmark in the non-reactive open-loop setting in Table 2, and evaluate autonomous driving models Chitta2023PAMI ; liao2024diffusiondrive trained solely on real data in both the real open-loop environment and its simulated counterpart in RealEngine, comparing the performance differences.
We use the the Predictive Driver Model Score (PDMS) Dauner2024NEURIPS between trajectories as a measure of similarity, and conduct evaluations only on sequences where the PDMS is greater than zero. The gap is defined as follows:
@@ Appendix B Consistency of driving behaviors
Specifically, $\mathrm{PDMS}_{\mathrm{real}}$ refers to the PDMS of the trajectory planned based on real sensor inputs, whereas $\mathrm{PDMS}_{\mathrm{sim}}$ corresponds to that planned using sensor data simulated by RealEngine. To evaluate the realism of camera and LiDAR simulation independently, we test three settings: (i) simulating only the camera ($\mathrm{PDMS}_{\mathrm{cam}}$), (ii) simulating only the LiDAR ($\mathrm{PDMS}_{\mathrm{lidar}}$), and (iii) simulating both camera and LiDAR ($\mathrm{PDMS}_{\mathrm{both}}$), as shown in Table 7. To better understand the realism of RealEngine, we also visualize the trajectories planned by DiffusionDrive liao2024diffusiondrive , as shown in Figure 11.
Table 7: Consistency of driving behaviors. We test Transfuser Chitta2023PAMI and DiffusionDrive liao2024diffusiondrive , which take camera and LiDAR inputs, under the three settings described in Appendix B.
@@
TransFuser Chitta2023PAMI 81.18 80.99 0.77% 80.78 0.55% 80.56 0.98%
DiffusionDrive liao2024diffusiondrive 80.75 81.07 0.89% 80.85 0.42% 81.07 1.07%
Figure 11: Consistency of driving behaviors. We test DiffusionDrive liao2024diffusiondrive under the same settings in Table 7. The close alignment between trajectories in the real and simulated environments demonstrates the high fidelity of our RealEngine.
@@
Appendix C Limitations
We still have a few limitations for future improvement:
Firstly, our current foreground meshes rely on existing datasets reizenstein21co3d ; du20243drealcar and 3D asset libraries, which limits the diversity of the foreground. We plan to incorporate generative models to generate foreground objects, thereby reducing manual effort, enhancing the richness of the scenes, and improving the generalizability of the simulation platform.
Secondly, although we have relighting models, there are still noticeable differences between the inserted traffic participants and the background scene. To address this issue, we plan to continuously track and integrate state-of-the-art advancements in foreground insertion to ensure seamless blending of traffic participants with the background.
Thirdly, we employ a simplified vehicle model that does not account for aerodynamic effects, suspension, or tire characteristics. Additionally, we overlook substantial variations in road conditions, such as changes in friction coefficients (e.g., wet or icy surfaces), gradients, and irregularities.

### v2

@@ A.1.3 LiDAR reconstruction
GS-LiDAR (Jiang et al., 2025) 0.315 6.930 21.25 0.0903 20.88
A.2 Consistency of driving behaviors
We replicate the same real-world evaluation scenarios from the Navsim (Dauner et al., 2024) benchmark in the non-reactive open-loop setting in Table˜2, and evaluate autonomous driving models (Chitta et al., 2023; Liao et al., 2025) trained solely on real data in both the real open-loop environment and its simulated counterpart in RealEngine, comparing the performance differences.
We use the the Predictive Driver Model Score (PDMS) (Dauner et al., 2024) between trajectories as a measure of similarity, and conduct evaluations only on sequences where the PDMS is greater than zero. The gap is defined as follows:
@@ A.2 Consistency of driving behaviors
Specifically, $\mathrm{PDMS}_{\mathrm{real}}$ refers to the PDMS of the trajectory planned based on real sensor inputs, whereas $\mathrm{PDMS}_{\mathrm{sim}}$ corresponds to that planned using sensor data simulated by RealEngine. To evaluate the realism of camera and LiDAR simulation independently, we test three settings: (i) simulating only the camera ($\mathrm{PDMS}_{\mathrm{cam}}$), (ii) simulating only the LiDAR ($\mathrm{PDMS}_{\mathrm{lidar}}$), and (iii) simulating both camera and LiDAR ($\mathrm{PDMS}_{\mathrm{both}}$), as shown in Table˜8. To better understand the realism of RealEngine, we also visualize the trajectories planned by DiffusionDrive (Liao et al., 2025), as shown in Figure˜11.
Table 8: Consistency of driving behaviors. We test Transfuser (Chitta et al., 2023) and DiffusionDrive (Liao et al., 2025), which take camera and LiDAR inputs, under the three settings described in Section˜A.2.
@@
TransFuser (Chitta et al., 2023) 81.18 80.99 0.77% 80.78 0.55% 80.56 0.98%
DiffusionDrive (Liao et al., 2025) 80.75 81.07 0.89% 80.85 0.42% 81.07 1.07%
Figure 11: Consistency of driving behaviors. We test DiffusionDrive (Liao et al., 2025) under the same settings in Table˜8. The close alignment between trajectories in the real and simulated environments demonstrates the high fidelity of our RealEngine.
@@
Moreover, to provide a quantitative downstream evaluation, we test the performance of the perception module of DiffusionDrive (Liao et al., 2025) with respect to the ground-truth foreground inserted. Specifically, we measuring the IoU between the model predicted bounding boxes and the ground-truth annotations on the original dataset, as well as the IoU between the model predicted bounding boxes and the inserted foreground annotations for the inserted objects. A smaller gap between the two IoUs indicates that the inserted foreground objects are more realistic and better perceived by the autonomous driving model.
As shown in Table˜9, although our relighting is not yet perfect, the inserted foreground objects can still be effectively perceived by the autonomous driving model. Moreover, the model’s ability to react to and avoid the inserted objects in certain simple scenarios further supports the realism of our insertions.
Table 9: Consistency of perception module. We evaluate the Intersection over Union (IoU) between the model-predicted bounding boxes of DiffusionDrive (Liao et al., 2025) and the ground-truth annotations in the original dataset, as well as the IoU between the model predictions and the annotations of the inserted foreground objects. A smaller discrepancy between these two IoU values suggests that the inserted objects are more realistic and are better recognized by the autonomous driving model.
[…]
DiffusionDrive (Liao et al., 2025) 72.3 68.9 4.7%